OpenStudy (anonymous):

Which ratio forms a portion With 9/15 A.3/6 B.2/3 C.12/30 D.6/10

OpenStudy (welshfella):

Divide 9 by 15 to give a decimal value Which of the options gives the same result?

OpenStudy (welshfella):

Using my calculator I get 9/15 = to 0.6

OpenStudy (anonymous):

Is it d or a

OpenStudy (anonymous):

You're basically trying to find \[\large \sf \frac{9}{15} = \frac{a}{b}\] So the easiest way is to cross multiply, so \[\large \sf 9b=15a\] For the first option, using this method, it would go like this \[\large \sf \frac{9}{15}=\frac{3}{6}\]\[\large \sf 9 \times 6 = 15 \times 3\]\[\large \sf 54 \neq 45\] so that isn't the right option, and you can do the rest

OpenStudy (mathmale):

I believe you're dealing with a"proportion," not a "portion." Check that out, please. Hint: reduce the fraction 9/15 as much as possible.

OpenStudy (anonymous):

Is it 3/6

OpenStudy (mathstudent55):

A proportion is two ratios that are equal. A ratio can be written in different ways, such as 2 to 3, 2:3, and 2/3, for example. If you write the ratio as a fraction, then a proportion is writing 2 equal fractions.

OpenStudy (mathstudent55):

You are given one ratio as a fraction: 9/15 Since you need to write a proportion, you need another fraction that is equivalent to 9/15. You need to pick a choice that has a fraction equivalent to 9/15. These are the steps to find your answer. 1. Reduce the given fraction 9/15 completely. 2. Reduce each choice completely until you find a choice that is equal to the reduced given fraction you found in step 1.

OpenStudy (mathstudent55):

Can you reduce 9/15? To reduce a fraction, divide the numerator and denominator by the largest number they are both divisible by. What number are 9 and 15 divisible by?

OpenStudy (mathmale):

Can you divide both 9 and 15 by 9 with no remainder? by 7? by 6? by 4? by 3? and so on.

jhonyy9 (jhonyy9):

so you know that 9/15 = (3*3)/(3*5) = ? what can you simplify there ? and now from these choices where you will get same result after you have simplified them A.3/6 = 3/(2*3) = ? B.2/3 C.12/30 = (2*2*3)/(2*3*5)= ? D.6/10 = (2*3)/(2*5) = ? hope these will help you